An exam worth 145 points contains 50 questions.Some of the questions are worth two points and some are worth five points. How many two point questions are on the test? How many five point questions are on the test? Write a linear system of equations that can be used to solve this problem. Solve
Answer by richwmiller(17219)

x+y=50
x=50-y
2*x+5*y=145
2*(50-y)+5*y=145
solve for y
100-2y+5y=145
3y=45
y=15
x=50-15
x=35
check
2*35+5*15=145
70+75=145
145=145
ok